Translate the problem into a pair of linear equations in two vaiables
How many liters of a 40%-alcohol solution must be mixed with 10 liters of a solution that is 80% alcohol to get a solution that is 60% alcohol

Answer by checkley77(12844)

.8x+.4*10=.6(10+x)
.8x+4=6+.6x
.8x-.6x=6-4
.2x=2
x=2/.2
x=10 liters of the 80% solution is used.
Proof:
.8*10+.4*10=.6(10+10)
8+4=.6*20
12=12